The dimension of a closed rctanqular box are 3ft by 6ft by 4ft. find the box's a. total surface area b. volume
Answer by vihits13(20)

surface area= 3*6*2 + 3*4*2 + 6*4*2= 36+24+48= 108 square feet.
Volume= 3*6*4= 72 cubic feet.